# LightRetriever: A LLM-based Hybrid Retrieval Architecture with 1000× Faster Query Inference

### v2

@@ 3.2 Main results
E5-Mistral 56.9 62.1 68.9 73.6 61.8 75.1 81.6 87.1
Llama3.2-1b
Baseline 53.1 60.2 66.9 72.0 63.8 78.1 84.1 87.9
Light Dense 48.7 55.8 63.0 68.3 58.7 73.1 79.7 83.8
Light Sparse 47.6 54.5 61.5 66.7 53.9 68.8 75.7 80.1
Light Hybrid 52.0-${}_{\text{-1.1}}$ 58.1 65.1 70.1 60.8-${}_{\text{-3.0}}$ 74.8 81.2 85.0
Llama3.2-3b
Baseline 55.6 62.9 69.7 74.7 66.1 81.0 86.6 90.0
Light Dense 51.0 58.6 65.5 70.5 60.4 75.1 81.5 85.3
Light Sparse 48.4 55.5 62.7 67.7 53.0 68.0 75.0 79.6
Light Hybrid 53.5-${}_{\text{-2.1}}$ 59.9 66.8 71.7 61.7-${}_{\text{-4.4}}$ 76.3 82.3 86.3
Llama3.1-8b
Baseline 56.8 64.3 71.1 75.7 67.6 82.0 87.4 90.9
Light Dense 52.0 59.5 66.5 71.4 61.4 75.6 82.3 86.4
Light Sparse 49.5 56.9 64.1 69.1 54.8 69.8 77.3 81.8
Light Hybrid 54.4-${}_{\text{-2.4}}$ 60.9 67.7 72.8 63.0-${}_{\text{-4.6}}$ 77.3 83.6 87.4
Mistral0.3-7b
Baseline 57.5 64.1 71.0 75.7 65.9 80.3 86.5 90.2
Light Dense 52.7 59.2 66.5 71.3 57.4 72.1 79.0 83.6
Light Sparse 48.0 56.1 63.3 68.3 50.7 65.5 72.7 77.8
Light Hybrid 54.5-${}_{\text{-3.0}}$ 60.1 67.4 72.1 59.4-${}_{\text{-6.5}}$ 73.9 80.6 84.8
Qwen2.5-1.5b
Baseline 54.0 60.5 67.5 72.3 65.9 81.0 86.7 90.4
Light Dense 48.9 56.0 63.1 68.1 61.9 77.0 83.2 87.4
Light Sparse 47.3 54.3 61.4 66.4 52.5 67.4 74.9 79.9
Light Hybrid 52.1-${}_{\text{-1.9}}$ 58.2 65.1 70.0 63.8-${}_{\text{-2.1}}$ 78.3 84.5 88.3
Qwen2.5-3b
Baseline 54.9 62.0 68.7 73.5 69.4 83.8 88.9 91.8
Light Dense 50.1 57.2 64.2 69.3 64.7 79.4 85.2 88.8
Light Sparse 47.5 54.9 62.0 67.2 54.1 69.2 76.5 81.1
Light Hybrid 52.8-${}_{\text{-2.1}}$ 59.1 66.0 70.9 65.7-${}_{\text{-3.7}}$ 80.2 86.1 89.3
Qwen2.5-7b
Baseline 56.6 63.6 70.5 75.2 70.1 84.4 89.4 92.4
Light Dense 50.6 58.6 65.6 70.5 65.0 79.8 85.8 89.3
Light Sparse 48.2 56.0 63.5 68.5 55.0 69.7 77.0 82.1
Light Hybrid 53.8-${}_{\text{-2.8}}$ 60.3 67.5 72.5 66.5-${}_{\text{-3.6}}$ 81.1 86.7 89.7
Retrieval performances
The retrieval performances are presented in Table 2. Experiments are conducted with 3 series of LLMs with different sizes. For each base model, we test the full-sized baseline and dense/sparse/hybrid (dense + sparse) retrieval performances of LightRetriever (Light). As is shown in the results, the performance gaps between LightRetriever (Hybrid) and its corresponding baseline are averaged around 5%, ranging from 2.1% (-1.1 point) to 9.9% (-6.5 point). Mistral [26] gets the highest nDCG@10 scores on the English BeIR benchmark, while Llama models [67] are better at recalls of BeIR. Qwen series [2] performs the best on the Chinese CMTEB Retrieval benchmark, attributed to better Chinese pre-training of Qwen.
3.3 Controls of vector dimension and sparsity
Retrieval systems usually involve indexing millions of documents into search databases, which consumes large amounts of disk space, especially for dense vectors. For example, indexing 8.8M MS-MARCO passages [52] with Llama-8b model consumes 134.92 GB for Faiss dense retrieval and 11.24 GB for Lucene sparse retrieval. Thus, it’s essential to explore the possibility of controlling index sizes by shrinking vectors. 1) For dense retrieval, we incorporate Matryoshka Representation Learning (MRL) [35] with LightRetriever to support flexible controls of dense vector dimensions. Given a dense vector $v^{den}$ with dimension $H$, MRL derives dimension-shrunk vectors by simply cutting out the Top-k dimensions $v^{den}[:k]$, then applies the same training loss as the original vectors. MRL trains a collection of Top-k dimensions together, enabling flexible choosing of a proper k. 2) For sparse retrieval, the sparse vector directly learns the term impact (i.e., lexical frequency) within a large sparse dimension. Thus, LightRetriever supports controlling sparsity by directly taking the Top-k terms, where the larger impact terms matter more.
Benchmark w/ KL wo/ KL
BeIR CMTEB-R BeIR CMTEB-R
Baseline 56.8 67.6 57.0+0.2 67.9+0.3
Light Dense 52.0 61.4 50.8-1.2 60.1-1.3
Light Sparse 49.5 54.8 48.7-0.8 54.4-0.4
Light Hybrid 54.4 63.0 53.9-0.5 62.1-0.9
3.4 Effect of auxiliary KL loss
Previous works for BERT-based asymmetric retrievers [73, 32] emphasize the importance of alignment between asymmetric students and full symmetric teachers. Our work also uses the KL loss as an auxiliary. Results in Table 3 show that removing such loss incurs limited drops of -0.5 and -0.9 on BeIR and CMTEB-R. Thus, such auxiliary KL loss is not mandatory in our experiment settings.
Table 4: Ablations of dense and sparse retrieval for LightRetriever (Llama-3.2-1b).
Dense Ablation BeIR CMTEB-R
LightRetriever 48.7 58.7
Symmetry
w/ Qry Full - Doc Light 8.7-40.0 9.4-49.3
w/ Qry Light - Doc Light 34.9-13.8 40.1-18.6
Other Ablations
w/ Direct Emb Layer 37.5-11.2 41.3-17.4
wo/ Instruction 47.1-1.6 56.1-2.6
Sparse Ablation BeIR CMTEB-R
LightRetriever 47.6 53.9
Symmetry
w/ Qry Full - Doc Light 16.2-31.4 10.8-43.1
Other Ablations
w/ Value fixed to 1 47.6+0.0 54.7+0.8
@@ 4 Ablations
4.1 Symmetry
Our work proposes an asymmetric architecture with an extremely lightweight query (Qry) encoder and a full-sized document (Doc) encoder. Besides, we also perform ablations on other symmetry settings, including lightweight on both query and document, and a full-sized query encoder with a lightweight passage encoder. Results in Table 4 show that a full-sized query encoder with a lightweight passage encoder leads to drastic performance degenerations, maybe caused by embedding collapses. Lightweight on both sides for dense retrieval works, but it does not perform excellently in retrieval tasks, just like Static Embedding [1]. Lightweight on both sides is not suitable for our sparse retrieval, because no learnable parameters are available.
4.2 Other ablations
Do we need a full query model during the dense training?
Dense retrieval of LightRetriever trains token vectors with full Transformers. Fast retrieval is achieved by caching learned dense token vectors and lookups during serving. If we directly use the Embedding Layer within the Transformers, the performances drop 11.2 and 17.4 points on BeIR and CMTEB-R. This verifies that during the training phase, a full-sized model is still needed for learning query token vectors.
Task instructions
Task instructions contribute +1.6 and +2.6 points on BeIR and CMTEB-R of dense retrieval. Similar effects of instructions are also observed in previous LLM retrievers [71, 5].
Value of sparse query vector
The sparse query vector is a key-value type of mapping. In Equation 7, we model the value of sparse query vectors by counting the corresponding token ids that appear in the query. For ablations, we choose a simpler way of modeling by stopping the counting and fixing the value to 1. In Table 4, BeIR results do not change. However, CMTEB-R results improve by +0.8 points. This indicates that counting term frequency in our sparse query vector seems unimportant to some extent, under large amounts of training on LightRetriever. The co-occurrence is more important.
@@ A.2.1 Text retrieval
Text retrieval builds on top of the vector space model [62], which encodes query and document representation vectors, and performs relevance search based on vector similarities.
@@
Dense retrieval [29] trains PLM-based encoders to condense queries and passages into low-dimensional dense vectors, then performs relevance search based on the Maximum Inner Product Search (MIPS) [51] algorithm. In recent years, dense retrievers, such as BGE [11] and E5 [70], have gained popularity for strong retrieval abilities across different tasks [50, 65], languages [83], and context granulaties [74, 52], owing to diverse training data [59], retrieval-customized pre-training [21, 75, 46], improved negative-mining [78], and so on. Recent SOTA retrievers [71, 5, 38] start to utilize LLMs [26, 67] as backbone encoders. With larger parameters and pre-training data, LLM-based retrievers enable significantly powerful retrieval capacities and domain adaptation abilities. However, LLMs also incur heavy online workloads when serving as query encoders. Improving the serving efficiency of LLM-based retrievers is still left for exploration.
@@
Traditional statistical sparse retrieval algorithms (e.g., TF-IDF [63], and BM25 [60]) do not involve learnable language models. Instead, they directly analyze lexical frequencies (e.g., TF) and importance (e.g., IDF) in a large vocabulary space. These lexicals are sparsely distributed and require building inverted indexes [45] for feasible similarity metric calculation, such as Apache Lucene [6] and Tantivy [58]. In recent years, sparse retrieval has started to work with pre-trained LMs (PLMs). BGE [11], SPLADE [20], SparTerm [3], and SPARTA [84] use customized XLM-RoBERTa [14] or BERT [18] as backbone encoders to learn term frequencies in an end-to-end manner and directly perform impact search based on the dot product of sparse vectors. Because of sparsity, these works also search with inverted indexes. Recent works [85] also try to incorporate LLMs with sparse retrieval. However, the online efficiency issue still exists.
@@
Hybrid retrieval is a technique that combines retrieval scores [72] or ranks [15] from multiple retrieval systems, such as dense [29] / sparse [60, 20] retrieval, and single [29] / multiple [31, 11] vector retrieval. It enables better retrieval abilities by interpolating results from multiple sources. Linear interpolation [72] of dense and sparse scores is used in our work.
@@ A.2.2 Improving inference efficiency
Previous works have tried to distill small query encoders from large encoders for better query inference efficiency. KALE [9] prunes the BERT-based query encoder and distills it during post-training. However, when the query encoder is pruned to one Transformers layer, its recall@20 on NQ [36] decreases by 25-30%. [73] proposes asymmetric dual-encoders by pruning BERT layers or using smaller BERT models. Its experiments show that a 2-layer BERT query encoder retains 92.5% of the full-sized performances via distillation and proper student initialization. USTAD [32] also proposes similar asymmetry and distillation strategies, which choose the 12-layer BERT as teachers and the 6-layer DistilBERT or 2-layer BERT-mini as students. Its experiments show that the distilled BERT-mini query encoder achieves 95-97% of the teacher’s performance because of better distillation.
However, these studies have primarily focused on small language models (SLMs), such as BERT [18], and overlooked feasible approaches for improving online inference efficiency of LLM-based retrievers, where such optimization is even more critical. Furthermore, these existing methods typically retain Transformer layers in the query encoders, which limits their efficiency gains. In addition, most of these works rely on SLMs and are trained on relatively narrow datasets, such as MS-MARCO [52], thus unable to compete with LLM-based retrievers.
@@
Previous works have also been dedicated to reducing index sizes for smaller disk space usage. Matryoshka representation learning (MRL) [35] supports adjusting the dense vector dimensions by directly taking the top-k dimension of the original vector as shrink vectors. It aligns and trains the corresponding multi-dimensional vectors in the end-to-end contrastive learning. For example, if $k\in\{64,128,256\}$, query-passage dense vectors with $\{64,128,256\}$-dims could be obtained together after one encoder forward operation. For LM-based sparse retrieval, the similarity scores are directly affected by the impacts of corresponding lexicals [20]. Thus, it naturally supports reducing the index sizes by retaining Top-k terms/lexicals. To explore the potential of controlling index sizes, our work also explores shrinking index sizes with dense MRL and sparse Top-k in the Experiments Section 3.3, which also shows strong generalization capacities.
A.3 Training datasets

### v3

@@ 3.1 Settings
[…]
E5-Mistral7b 56.9 62.1 68.9 73.6 61.8 75.1 81.6 87.1
Full Symmetric Retrievers
Llama3.2-1b 53.1 60.2 66.9 72.0 63.8 78.1 84.1 87.9
Llama3.2-3b 55.6 62.9 69.7 74.7 66.1 81.0 86.6 90.0
Llama3.1-8b 56.8 64.3 71.1 75.7 67.6 82.0 87.4 90.9
Qwen2.5-1.5b 54.0 60.5 67.5 72.3 65.9 81.0 86.7 90.4
Qwen2.5-3b 54.9 62.0 68.7 73.5 69.4 83.8 88.9 91.8
Qwen2.5-7b 56.6 63.6 70.5 75.2 70.1 84.4 89.4 92.4
LightRetriever
Llama3.2-1b 52.0-$_{\text{-1.1}}$ 58.1 65.1 70.1 60.8-$_{\text{-3.0}}$ 74.8 81.2 85.0
Llama3.2-3b 53.5-$_{\text{-2.1}}$ 59.9 66.8 71.7 61.7-$_{\text{-4.4}}$ 76.3 82.3 86.3
Llama3.1-8b 54.4-$_{\text{-2.4}}$ 60.9 67.7 72.8 63.0-$_{\text{-4.6}}$ 77.3 83.6 87.4
Qwen2.5-1.5b 52.1-$_{\text{-1.9}}$ 58.2 65.1 70.0 63.8-$_{\text{-2.1}}$ 78.3 84.5 88.3
Qwen2.5-3b 52.8-$_{\text{-2.1}}$ 59.1 66.0 70.9 65.7-$_{\text{-3.7}}$ 80.2 86.1 89.3
Qwen2.5-7b 53.8-$_{\text{-2.8}}$ 60.3 67.5 72.5 66.5-$_{\text{-3.6}}$ 81.1 86.7 89.7
Performance comparisons
[…]
Notably, Static Embedding performs retrieval using a single shared embedding layer for both queries and documents, offering low retrieval performance but highly efficient baseline. BM25, as a classical term-frequency-based method, remains a widely adopted sparse retriever. The combination of them thus provides a reference for efficient retrieval on both sides. In contrast, BGE-m3 is a Transformer-based retriever trained with multi-stage pretraining and fine-tuning, and represents state-of-the-art performance in hybrid retrieval.
Except for Static Embedding and BM25, all compared retrievers rely on deep Transformer encoders on both query and document sides. LightRetriever is, to our knowledge, the first method to address the online efficiency bottleneck on the query side of LLM-based retrievers, striking a balance between retrieval quality and inference efficiency.
3.2 Main results
Encoding speeds
Table 1 presents a detailed breakdown of retrieval time for processing 65,536 Bing queries over 1 million MS MARCO passages, highlighting the efficiency advantages of LightRetriever over full symmetric retrievers. Among all time consumptions, model encoding is the dominant contributor to latency in LLM-based retrievers. For example, Full-Llama8b and Full-Qwen7b require over 100 seconds solely for query encoding (109.5s and 100.7s, respectively), making them a bottleneck for large-scale, real-time retrieval.
In contrast, LightRetriever reduces encoding time to below 50 ms (0.0412–0.0420s), achieving over a 2500× speedup in the encoding phase. This drastic reduction is made possible by replacing full forward passes of large Transformer models with a simple embedding lookup. And the overall throughput achieves over 10x QPS speedup. Notably, this lightweight encoding does not significantly compromise retrieval performance. LightRetriever maintains competitive nDCG@10 scores on both English (BeIR) and Chinese (CMTEB-Retrieval) benchmarks.
These results prove the effectiveness of LightRetriever’s decoupled query modeling paradigm, which drastically improves online query encoding efficiency while preserving good retrieval quality, making it suitable for latency-critical applications.
Performance comparisons
Table 2 compares retrieval effectiveness across a wide range of retrievers on both English (BeIR) and Chinese (CMTEB-Retrieval) benchmarks. The main focus of this comparison is to evaluate the performance degradation introduced by LightRetriever’s lightweight query encoding. Overall, by leveraging deep Transformer encoders on both query and document sides, full symmetric retrievers achieve the highest retrieval scores. For instance, Llama3.1-8b performs best on BeIR (56.8 nDCG@10), while Qwen2.5-7b reaches the best performance on CMTEB-R (70.1 nDCG@10). However, such performance comes at the cost of significant computational overhead, as discussed above.
[…]
These results demonstrate that LightRetriever effectively balances retrieval quality and efficiency. By avoiding expensive query-side encoding and maintaining high retrieval performance, it establishes a promising new paradigm for scalable LLM-based retrieval systems, especially in latency-sensitive or high-throughput scenarios.
@@ 4 Ablations
Our core design adopts an asymmetric architecture, with a lightweight query MLP Embedding and a full-sized document model for inference. Specifically, we utilize a full-sized model on both sides at training time, then cache all query token embeddings in one Embedding Lookup for fast query-side inference. In ablation studies below (A1-A3), we are interested in other potential symmetric/asymmetric architectures, which are described in the following Table 3.
Table 3: Symmetry Ablation settings (A1-A3).
Train Inference
Query Doc Query Doc
Ours Full Full MLP Full
A1 Full Full MLP MLP
A2 MLP Full MLP Full
A3 1-layer Transformers Full 1-layer Transformers Full
A1. Is symmetric lightweight inference effective?
To test the necessity of the above asymmetry, we evaluate a symmetric lightweight setup where both sides use simplified dense encoders. As shown in Table 4, this leads to severe performance drops on both BeIR (–13.8) and CMTEB-R (–18.6), confirming that full document representations are crucial for maintaining retrieval effectiveness. This symmetric setting is not functional for sparse retrieval due to the absence of learnable parameters.
The best symmetric lightweight retriever is Static Embedding (one MLP) + BM25 (term-based) in Table 2, with nDCG@10 of 44.7 on BeIR and 52.1 on CMTEB-R. However, due to the lack of deep modeling on documents, such a combination is still outperformed.
We also tested a reversed asymmetry with deep encoding for queries and lightweight documents. However, we observed training instability and degraded performance, likely caused by embedding collapse.
A2. Can we use only a query MLP for training?
To confirm that a deep query encoder is still needed during training, we replaced the Transformer-based query encoder with a simple MLP that reuses only the LLM’s embedding layer. This setup significantly degrades performance (–11.2 on BeIR, –17.4 on CMTEB-R), indicating that the full-sized modeling for query side training remains essential during training to learn effective token-level representations.
Table 4: Ablations results of LightRetriever
Ablations / nDCG@10 BeIR CMTEB-R
Llama-1b 48.7 58.7
w/ Both-side Lightweight 34.9-13.8 40.1-18.6
w/ Encq use a MLP 37.5-11.2 41.3-17.4
Llama-8b 54.4 63.0
w/ 1st-layer on query 50.1-4.3 54.4-8.6
A3. Is adding one Transformers layer better on the query side?
We also evaluated using only the first Transformer layer on the query side. While this reduces model size during training, it still underperforms the LightRetriever (–4.3 on BeIR, –8.6 on CMTEB-R) and introduces inference-time overhead, making it incompatible with the design goals of LightRetriever. This again proves the need for full-sized query modeling during the training phase.
Figure 5: Ablations of retrieval performances (nDCG@10) of LightRetriever on BeIR and CMTEB-R benchmarks.
A4. Breakdown Ablation of Retrieval Performance
Figure 5 reports LightRetriever’s performance under dense, sparse, and hybrid configurations. Dense retrieval consistently outperforms sparse across all backbones (e.g., LLaMA3, Qwen2.5) and model sizes (1B–8B).
While sparse retrieval alone is less effective, the hybrid approach recovers most of the performance gap with minimal overhead. For instance, LLaMA3-1B achieves 52.0 and 60.8 nDCG@10 on BeIR and CMTEB-R in the hybrid setting, with only 1.1 and 3.0 points below the full model, but with a much smaller query encoder. These results confirm the complementarity of both signals and the effectiveness of LightRetriever in balancing efficiency and accuracy.
@@ A.2.1 Text retrieval
Text retrieval builds on top of the vector space model Salton [1975], which encodes query and document representation vectors, and performs relevance search based on vector similarities.
@@
Dense retrieval Karpukhin et al. [2020] trains PLM-based encoders to condense queries and passages into low-dimensional dense vectors, then performs relevance search based on the Maximum Inner Product Search (MIPS) Mussmann and Ermon [2016] algorithm. In recent years, dense retrievers, such as BGE Chen et al. [2024] and E5 Wang et al. [2024a], have gained popularity for strong retrieval abilities across different tasks Muennighoff et al. [2023], Thakur et al. [2021], languages Zhang et al. [2023], and context granulaties Williams et al. [2018], Nguyen et al. [2016], owing to diverse training data Reimers [2019], retrieval-customized pre-training Gao and Callan [2021], Xiao et al. [2022], Ma et al. [2024], improved negative-mining Xiong et al. [2021], and so on. Recent SOTA retrievers Wang et al. [2024b], BehnamGhader et al. [2024], Lee et al. [2024] start to utilize LLMs Jiang et al. [2023], Touvron et al. [2023] as backbone encoders. With larger parameters and pre-training data, LLM-based retrievers enable significantly powerful retrieval capacities and domain adaptation abilities. However, LLMs also incur heavy online workloads when serving as query encoders. Improving the serving efficiency of LLM-based retrievers is still left for exploration.
@@
Traditional statistical sparse retrieval algorithms (e.g., TF-IDF Salton and Buckley [1988], and BM25 Robertson et al. [1994]) do not involve learnable language models. Instead, they directly analyze lexical frequencies (e.g., TF) and importance (e.g., IDF) in a large vocabulary space. These lexicals are sparsely distributed and require building inverted indexes Luhn [1957] for feasible similarity metric calculation, such as Apache Lucene Bialecki et al. [2012] and Tantivy Quickwit Inc. [2024]. In recent years, sparse retrieval has started to work with pre-trained LMs (PLMs). BGE Chen et al. [2024], SPLADE Formal et al. [2021], SparTerm Bai et al. [2020], and SPARTA Zhao et al. [2021] use customized XLM-RoBERTa Conneau et al. [2020] or BERT Devlin et al. [2019] as backbone encoders to learn term frequencies in an end-to-end manner and directly perform impact search based on the dot product of sparse vectors. Because of sparsity, these works also search with inverted indexes. Recent works Zhuang et al. [2024] also try to incorporate LLMs with sparse retrieval. However, the online efficiency issue still exists.
@@
Hybrid retrieval is a technique that combines retrieval scores Wang et al. [2021] or ranks Cormack et al. [2009] from multiple retrieval systems, such as dense Karpukhin et al. [2020] / sparse Robertson et al. [1994], Formal et al. [2021] retrieval, and single Karpukhin et al. [2020] / multiple Khattab and Zaharia [2020], Chen et al. [2024] vector retrieval. It enables better retrieval abilities by interpolating results from multiple sources. Linear interpolation Wang et al. [2021] of dense and sparse scores is used in our work.
@@ A.2.2 Improving inference efficiency
Previous works have tried to distill small query encoders from large encoders for better query inference efficiency. KALE Campos et al. [2023] prunes the BERT-based query encoder and distills it during post-training. However, when the query encoder is pruned to one Transformers layer, its recall@20 on NQ Kwiatkowski et al. [2019] decreases by 25-30%. Wang and Lyu [2023] proposes asymmetric dual-encoders by pruning BERT layers or using smaller BERT models. Its experiments show that a 2-layer BERT query encoder retains 92.5% of the full-sized performances via distillation and proper student initialization. USTAD Kim et al. [2024] also proposes similar asymmetry and distillation strategies, which choose the 12-layer BERT as teachers and the 6-layer DistilBERT or 2-layer BERT-mini as students. Its experiments show that the distilled BERT-mini query encoder achieves 95-97% of the teacher’s performance because of better distillation.
However, these studies have primarily focused on small language models (SLMs), such as BERT Devlin et al. [2019], and overlooked feasible approaches for improving online inference efficiency of LLM-based retrievers, where such optimization is even more critical. Furthermore, these existing methods typically retain Transformer layers in the query encoders, which limits their efficiency gains. In addition, most of these works rely on SLMs and are trained on relatively narrow datasets, such as MS-MARCO Nguyen et al. [2016], thus unable to compete with LLM-based retrievers.
@@
Previous works have also been dedicated to reducing index sizes for smaller disk space usage. Matryoshka representation learning (MRL) Kusupati et al. [2022] supports adjusting the dense vector dimensions by directly taking the top-k dimension of the original vector as shrink vectors. It aligns and trains the corresponding multi-dimensional vectors in the end-to-end contrastive learning. For example, if $k\in\{64,128,256\}$, query-passage dense vectors with $\{64,128,256\}$-dims could be obtained together after one encoder forward operation. For LM-based sparse retrieval, the similarity scores are directly affected by the impacts of corresponding lexicals Formal et al. [2021]. Thus, it naturally supports reducing the index sizes by retaining Top-k terms/lexicals. To explore the potential of controlling index sizes, our work also explores shrinking index sizes with dense MRL and sparse Top-k in the Section A.3, which also shows strong generalization capacities.
A.3 Additional results about controls of vector dimension and sparsity
Retrieval systems usually involve indexing millions of documents into search databases, which consumes large amounts of disk space, especially for dense vectors. For example, indexing 8.8M MS-MARCO passages Nguyen et al. [2016] with Llama-8b model consumes 134.92 GB for Faiss dense retrieval and 11.24 GB for Lucene sparse retrieval. Thus, it’s essential to explore the possibility of controlling index sizes by shrinking vectors. 1) For dense retrieval, we incorporate Matryoshka Representation Learning (MRL) Kusupati et al. [2022] with LightRetriever to support flexible controls of dense vector dimensions. Given a dense vector $v^{den}$ with dimension $H$, MRL derives dimension-shrunk vectors by simply cutting out the Top-k dimensions $v^{den}[:k]$, then applies the same training loss as the original vectors. MRL trains a collection of Top-k dimensions together, enabling flexible choosing of a proper k. 2) For sparse retrieval, the sparse vector directly learns the term impact (i.e., lexical frequency) within a large sparse dimension. Thus, LightRetriever supports controlling sparsity by directly taking the Top-k terms, where the larger impact terms matter more.
A.4 Additional ablation about the effect of auxiliary KL loss
Benchmark w/ KL wo/ KL
BeIR CMTEB-R BeIR CMTEB-R
Dense 52.0 61.4 50.8-1.2 60.1-1.3
Sparse 49.5 54.8 48.7-0.8 54.4-0.4
Hybrid 54.4 63.0 53.9-0.5 62.1-0.9
Previous works for BERT-based asymmetric retrievers Wang and Lyu [2023], Kim et al. [2024] emphasize the importance of alignment between asymmetric students and full symmetric teachers. To explore the effect of such alignment, besides contrastive loss as the main loss function, we also explore to adopt the KL loss as an auxiliary to align the asymmetric similarity scores $S^{student}=v_{q}^{den}\cdot v_{d}^{den}$ to full-sized symmetric similarity scores $S^{teacher}=Enc_{q}(q)\cdot Enc_{d}(d)$,
$\ell^{Align}=\texttt{KL-Div}(S^{student},S^{teacher})$ (9)
Results in Table 5 show that removing such loss incurs limited drops of -0.5 and -0.9 on BeIR and CMTEB-R. Thus, such auxiliary KL loss is not mandatory needed in our experiments.
A.5 Training datasets